Cycle Pths, Inc makes bicycles, trycicles and unicycles. Last week they made 88 more bicycles than unicycles, and 5 times as many tricycles as unicycles. If they made 40 more bicycles than tricycles, how many unicycles did they make?
Answer by HyperBrain(694)

Let x=the no. of bicycles
y=no.of tricycles
z=no. of unicycles
(1)x=z+88
(2)y=5z
(3)x=y+40
Substituting (1) and (2) to (3),
z+88=5z+40
88=4z+40
48=4z
z=12
Therefore, they made 12 unicycles!
HyperBrain